OpenStudy (animelover91804):

What is the value of the expression? –13.7 + (–5 + 2.7) A. –21.4 B. –16 C. –11.4 D. –6

OpenStudy (animelover91804):

Let's see who knows! i already know though XD

OpenStudy (alexandervonhumboldt2):

-13.7+(-5+2.7)=-13.7-5+2.7=-11-5=-14

OpenStudy (alexandervonhumboldt2):

woops -16

OpenStudy (anonymous):

just put it into a calculator, or google https://www.google.com/search?q= –13.7+%2B+(–5+%2B+2.7) its -16